# A study of bound states of $\eta,\eta^{{}^{\prime}},D^{0}$, $\bar{D^{0}}$, $B^{0}$, $\bar{B^{0}}$, $\bar{K^{0}}$ and $\phi$ mesons with light and heavy nuclei within chiral SU(3) model

### v1

@@ IV.3 Binding energies for $B^{0}$ and $\bar{B^{0}}$ mesic nuclei
The studies on the bound states of $B$ meson with nuclei has not been done much in literature [44]. In the present section we shall discuss the formation of bound states of pseudoscalar $B^{0}$ and $\bar{B^{0}}$ mesons by calculating their optical potential using the chiral SU(3) model. In Fig. 5, the mass shift of these pseudoscalar mesons are plotted as a function of baryon density $\rho_{B}$ (in units of $\rho_{0}$) for isospin asymmetry $I=0,0.211$ and $0.3$. Both pseudoscalar $B^{0}$ and $\bar{B^{0}}$ mesons are observed to undergo appreciable mass drop at nuclear matter density. However, the isospin asymmetry effects are not appreciable. In symmetric nuclear medium the values of $\Delta m$ for $B^{0}$ and $\bar{B^{0}}$ mesons are observed to be $-48.69$ and $-74.06$ MeV, whereas at $4\rho_{0}$ these values changes to $-356.40$ and $-453.28$ MeV, respectively. As shown in Fig. 6, for a given nuclei, at the center, the real part of optical potentials have larger value for $\bar{B^{0}}$ as compared to $B^{0}$ mesons. The values of binding energy and decay width for these pseudoscalar mesosn are given in Table 4 and 5. As expected from the magnitude of mass drop for $B^{0}$ and $\bar{B^{0}}$ mesons, the magnitude of binding energy with given nuclei is more for $\bar{B^{0}}$ mesons as compared to $B^{0}$.
@@
Figure 7: The mass shift of $\bar{K^{0}}$ mesons is shown as a function of density $\rho_{B}$ (in units of nuclear saturation density $\rho_{0}$) of nuclear medium for isospin asymmetry, $I=0,0.211$ and $0.3$.
@@
Figure 9: The mass shift [in subplot (a)] and decay width [in subplot (b)] of $\phi$ mesons are shown as a function of density $\rho_{B}$ (in units of nuclear saturation density $\rho_{0}$) of nuclear medium for isospin asymmetry, $I=0,0.211$ and $0.3$.
@@ IV.4 Binding energies for $\bar{K^{0}}$ and $\phi$ mesic nuclei
Lastly, in this subsection we explore the formation of bound states for $\bar{K^{0}}$ and $\phi$ mesons. Within the chiral SU(3) model, among the neutral $K^{0}$ and $\bar{K^{0}}$ mesons, only latter is observed to undergo mass drop as a function of density of the nuclear medium [65, 66, 64] and therefore, we consider $\bar{K^{0}}$ neutral meson only to explore the bound state formation. In Fig. 7 we have shown the variation of mass shift for $\bar{K^{0}}$ mesons as a function of baryon density $\rho_{B}$. The plot of real part of optical potential $U(r)$ as a function of distance $r$ from the center of the given nuclei of mass number $A$ is given for $\bar{K^{0}}$ meson in Fig. 8. At $\eta=0$, the values of mass shift for $\bar{K^{0}}$ mesons are found to be $-36.90$ and $-184.31$ MeV, at $\rho_{B}=\rho_{0}$ and $4\rho_{0}$, respectively. For a given density of medium, an increase in the isospin asymmetry enhances the mass drop for $\bar{K^{0}}$ mesons, which is more appreciable at higher baryon densities. As can be seen from Table 6, for the nuclei 12C, 18O and 40Ca, the negative values of binding energies are observed only for $1s$ state. Further, in this case when the value of $\kappa$ is considered finite, it may be difficult to observe the bound state of $\bar{K^{0}}$ meson with these nuclei due to comparable or larger value of decay width $\Gamma$.
@@
As we discussed earlier in Sec. III.5, the mass shift and decay width of $\phi$ meson are obtained in the present work by calculating the self energy, which involve the interactions of $\phi$ mesons with $K$ and $\bar{K}$, at one loop level. In Fig. 9 we have plotted the mass shift and decay width of $\phi$ mesons as a function of $\rho_{B}/\rho_{0}$. The values of mass shift $\Delta m_{\phi}^{*}$ and decay width $\ \Gamma_{\phi}^{*}$ at baryon density $\rho_{B}=\rho_{0}(4\rho_{0})$ are observed to be $-2.00(-20.28)$ and $4.96(27.68)$ MeV, respectively, for the value of cuttoff parameter, $\Lambda_{c}=3000$ MeV. In Fig. 10 the real and imaginary optical potentials, $U_{\phi}(r)$ and $W_{\phi}(r)$ of $\phi$ mesons are plotted as a function of distance $r$ for the four nuclei ($A=12,16,40$ and $208$). As we observed very small negative mass shift for the $\phi$ mesons, the bound states are not possible for these mesons within the present calculations where the in-medium masses of $K$ and $\bar{K}$ mesons are evaluated using the chiral SU(3) model. In the chiral SU(3) model, the properties of kaons and antikaons are modified differently in the nuclear medium. The effective mass of kaons increases whereas the mass of antikaons decrease as the density of the medium is changed from zero to finite value [65, 66, 64]. In Ref. [6], the $\phi$ meson bound states are investigated considering the modification of kaons and antikaons within the QMC model. The in-medium masses of kaons and antikaons are considered same in the nuclear medium and are observed to decrease as a function of baryon density $\rho_{B}$. Thus, the mass drop in the $\phi$ mesons is also observed to be large as compared to our present calculations. For example, at $\rho_{B}=\rho_{0}$, $\Delta m_{\phi}$ is observed to be $\approx 20$ MeV for the cuttoff $\Lambda_{c}=2000$ MeV. This larger mass drop in the $\phi$ meson masses also leads to the possibility of bound states in the calculations of Ref. [6].
@@ V Summary
For the study of bound states for pseudoscalar $\eta$ and $\eta^{{}^{\prime}}$ mesons within the chiral SU(3) model, we considered the mixing between the states $\eta_{8}$ and $\eta_{0}$. We observed appreciable mass drop for the $\eta$ mesons whereas $\eta^{{}^{\prime}}$ mass is observed to decrease very less. This leads to the possibility of bound states of $\eta$ mesons with the nuclei considered in this work, whereas for the $\eta^{{}^{\prime}}$ meson no bound state is observed. Using the chiral SU(3) model, generalized to SU(4) sector, we also elaborated on the formation of bound states for $D^{0}$ and $\bar{D^{0}}$ mesons. For both $D^{0}$ and $\bar{D^{0}}$ mesons the mass drop observed at nuclear saturation density is such that it lead to formation of bound states. Among $D^{0}$ and $\bar{D^{0}}$, former is observed to undergo larger mass drop and this lead to the formation of bound states with higher states also. For the case of $B^{0}$ and $\bar{B^{0}}$ mesons also the obtained negative mass shift lead to the formation of bound states for lower and excited states, especially for nuclei with larger mass number $A$. In our present work we also observed that the pseudoscalar $\bar{K^{0}}$ mesons can form the bound states as negative mass shift is observed within the chiral SU(3) model. For the case of vector $\phi$ mesons the mass drop obtained in current work is not large enough to form their bound states with the light and heavy nuclei. The present investigations on the formation of meson-nuclei bound states will be significant for the experiments of $\bar{\text{P}}$ at FAIR, WASA at COSY and J-PARC facilities.

### v2

@@ IV.3 Binding energies for $B^{0}$ and $\bar{B^{0}}$ mesic nuclei
The studies on the bound states of $B$ meson with nuclei has not been done much in literature [44]. In the present section we shall discuss the formation of bound states of pseudoscalar $B^{0}$ and $\bar{B^{0}}$ mesons by calculating their optical potential using the chiral SU(3) model. In Fig. 5, the mass shift of these pseudoscalar mesons are plotted as a function of baryon density $\rho_{B}$ (in units of $\rho_{0}$) for isospin asymmetry $I=0,0.211$ and $0.3$, at $T=0$ and $100$ MeV. Both pseudoscalar $B^{0}$ and $\bar{B^{0}}$ mesons are observed to undergo appreciable mass drop at nuclear matter density, whereas, for given density, an increase in temperature from zero to finite values causes an increase in the mass, i.e., less drop in mass is observed. However, the isospin asymmetry effects are not appreciable. In cold symmetric nuclear medium, $\rho_{B}=\rho_{0}$, the values of $\Delta m$ for $B^{0}$ and $\bar{B^{0}}$ mesons are observed to be $-48.69$ and $-74.06$ MeV, whereas at $4\rho_{0}$ these values changes to $-356.40$ and $-453.28$ MeV, respectively. As shown in Fig. 6, for a given nuclei, at the center, the real part of optical potentials have larger value for $\bar{B^{0}}$ as compared to $B^{0}$ mesons. In case of $B^{0}$ mesons the optical potentials have positive values for certain range whereas for $\bar{B^{0}}$ this is not the case. This is because of opposite contribution of Weinberg Tomozawa term to $B^{0}$ and $\bar{B^{0}}$ mesons (see first term of Eqs.(48) and (49)). For $B^{0}$ mesons, Weinberg Tomozawa term lead to positive values of optical potentials. The values of binding energy and decay width for these pseudoscalar mesons are given in Table 4 and 5. As expected from the magnitude of mass drop for $B^{0}$ and $\bar{B^{0}}$ mesons, the magnitude of binding energy with given nuclei is more for $\bar{B^{0}}$ mesons as compared to $B^{0}$.
@@
Figure 7: The mass shift of $\bar{K^{0}}$ mesons is shown as a function of density $\rho_{B}$ (in units of nuclear saturation density $\rho_{0}$) of nuclear medium for isospin asymmetry, $I=0,0.211$ and $0.3$. Results are shown for temperatures $T=0$ (subplot (a)) and $100$ MeV (subplot (b)).
@@
Figure 9: The mass shift [in subplots (a) and (c)] and decay width [in subplots (b) and (d)] of $\phi$ mesons are shown as a function of density $\rho_{B}$ (in units of nuclear saturation density $\rho_{0}$) of nuclear medium for isospin asymmetry, $I=0,0.211$ and $0.3$. Results are shown for temperatures $T=0$ and $100$ MeV.
@@
Figure 11: The average effective masses $m_{K}^{*}\left(=\frac{m_{K^{0}}^{*}+m_{K^{+}}^{*}}{2}\right)$ and $m_{\bar{K}}^{*}\left(=\frac{m_{\bar{K}^{0}}^{*}+m_{K^{-}}^{*}}{2}\right)$ for kaons and antikaons are plotted as a function of distance $r$ from center, for nuclei 12C and 208Pb, with mass number, $A=12$ and $208$, respectively.
@@ IV.4 Binding energies for $\bar{K^{0}}$ and $\phi$ mesic nuclei
Lastly, in this subsection we explore the formation of bound states for $\bar{K^{0}}$ and $\phi$ mesons. Within the chiral SU(3) model, among the neutral $K^{0}$ and $\bar{K^{0}}$ mesons, only latter is observed to undergo mass drop as a function of density of the nuclear medium [65, 66, 64] and therefore, we consider $\bar{K^{0}}$ neutral meson only to explore the bound state formation. In Fig. 7 we have shown the variation of mass shift for $\bar{K^{0}}$ mesons as a function of baryon density $\rho_{B}$, at temperatures $T=0$ and 100 MeV. The plot of real part of optical potential $U(r)$ as a function of distance $r$ from the center of the given nuclei of mass number $A$ is given for $\bar{K^{0}}$ meson in Fig. 8. At $I=0$ and $T=0$ (100) MeV, the values of mass shift for $\bar{K^{0}}$ mesons are found to be $-36.90(-36.43)$ and $-184.31(-175.87)$ MeV, at $\rho_{B}=\rho_{0}$ and $4\rho_{0}$, respectively. For a given density of medium, an increase in the isospin asymmetry enhances the mass drop for $\bar{K^{0}}$ mesons, which is more appreciable at higher baryon densities. As can be seen from Table 6, for the nuclei 12C, 18O and 40Ca, the negative values of binding energies are observed only for $1s$ state. Further, in this case when the value of $\kappa$ is considered finite, it may be difficult to observe the bound state of $\bar{K^{0}}$ meson with these nuclei due to comparable or larger value of decay width $\Gamma$.
@@
As we discussed earlier in Sec. III.5, the mass shift and decay width of $\phi$ meson are obtained in the present work by calculating the self energy, which involve the interactions of $\phi$ mesons with $K$ and $\bar{K}$, at one loop level. In Fig. 9 we have plotted the mass shift and decay width of $\phi$ mesons as a function of $\rho_{B}/\rho_{0}$. At temperature $T=0$ MeV, the values of mass shift $\Delta m_{\phi}^{*}$ and decay width $\ \Gamma_{\phi}^{*}$ at baryon density $\rho_{B}=\rho_{0}(4\rho_{0})$ are observed to be $-2.00(-20.28)$ and $4.96(27.68)$ MeV, respectively, for the value of cutoff parameter, $\Lambda_{c}=3000$ MeV. In Fig. 10 the real and imaginary optical potentials, $U_{\phi}(r)$ and $W_{\phi}(r)$ of $\phi$ mesons are plotted as a function of distance $r$ for the four nuclei ($A=12,16,40$ and $208$). As we observed very small negative mass shift for the $\phi$ mesons, the bound states are not possible for these mesons within the present calculations where the in-medium masses of $K$ and $\bar{K}$ mesons are evaluated using the chiral SU(3) model. In the chiral SU(3) model, the properties of kaons and antikaons are modified differently in the nuclear medium. The effective mass of kaons increases whereas the mass of antikaons decrease as the density of the medium is changed from zero to finite value [65, 66, 64]. From Fig. 10, we observe that the real part of the optical potential remains mostly repulsive. This is because of dominance of repulsive contributions of kaons as a function of $r$ over the attractive contribution of antikaons to the in-medium masses of $\phi$ mesons. As discussed in Sec. III.5, average masses $m_{K}^{*}\left(=\frac{m_{K^{0}}^{*}+m_{K^{+}}^{*}}{2}\right)$ and $m_{\bar{K}}^{*}\left(=\frac{m_{\bar{K}^{0}}^{*}+m_{K^{-}}^{*}}{2}\right)$ are used in the calculations of self-energy of $\phi$ mesons given by Eq. (54). As shown in Fig.11, average mass $m_{K}^{*}$ remains repulsive (in-medium mass greater than vacuum mass ($\sim$ 494 MeV)) as a function of $r$ whereas $m_{\bar{K}}^{*}$ is attractive (in-medium mass less than vacuum mass ($\sim$ 494 MeV)). For certain range of $r$ values, increased dominance of repulsive contribution causes a bump in real part of optical potential of $\phi$ mesons and a well in the imaginary part. In Ref. [6], the $\phi$ meson bound states are investigated considering the modification of kaons and antikaons within the QMC model. The in-medium masses of kaons and antikaons are considered same in the nuclear medium and are observed to decrease as a function of baryon density $\rho_{B}$. Thus, the mass drop in the $\phi$ mesons is also observed to be large as compared to our present calculations. For example, at $\rho_{B}=\rho_{0}$, $\Delta m_{\phi}$ is observed to be $\approx-20$ MeV for the cuttoff $\Lambda_{c}=2000$ MeV. This larger mass drop in the $\phi$ meson masses also leads to the possibility of bound states in the calculations of Ref. [6]. However, as discussed above, in the chiral SU(3) model considered in the present calculations, the in-medium mass of kaons increases whereas the mass of antikaon decreases as a function of density of nuclear medium. Actually, the effective masses of kaons and antikaons in the chiral SU(3) model are modified in the nuclear medium through the Weinberg Tomozawa term, scalar meson exchange term and range terms [65, 66, 64]. As a function of density of the nuclear medium, the leading order Weinberg Tomozawa term gives repulsive contributions to kaons and attractive for antikaons. Overall this leads to an increase in the mass of kaons and a decrease in the mass of antikaons as function of $\rho_{B}$. Since the in-medium masses of $\phi$ mesons are modified through the medium modifications of kaons and antikaons, their opposite behavior as a function of density leads to small mass shift in the mass of $\phi$ mesons in the present calculations.
@@ V Summary
For the study of bound states for pseudoscalar $\eta$ and $\eta^{{}^{\prime}}$ mesons within the chiral SU(3) model, we considered the mixing between the states $\eta_{8}$ and $\eta_{0}$. We observed appreciable mass drop for the $\eta$ mesons whereas $\eta^{{}^{\prime}}$ mass is observed to decrease very less. This leads to the possibility of bound states of $\eta$ mesons with the nuclei considered in this work, whereas for the $\eta^{{}^{\prime}}$ meson no bound state is observed. Using the chiral SU(3) model, generalized to SU(4) sector, we also elaborated on the formation of bound states for $D^{0}$ and $\bar{D^{0}}$ mesons. For both $D^{0}$ and $\bar{D^{0}}$ mesons the mass drop observed at nuclear saturation density is such that it lead to formation of bound states. Among $D^{0}$ and $\bar{D^{0}}$, former is observed to undergo larger mass drop and this lead to the formation of bound states with higher states also. For the case of $B^{0}$ and $\bar{B^{0}}$ mesons also the obtained negative mass shift lead to the formation of bound states for lower and excited states, especially for nuclei with larger mass number $A$. In our present work we also observed that the pseudoscalar $\bar{K^{0}}$ mesons can form the bound states as negative mass shift is observed within the chiral SU(3) model. For the case of vector $\phi$ mesons the mass drop obtained in current work is not large enough to form their bound states with the light and heavy nuclei. In Ref. [44], meson-nucleus bound states are studied using the QMC model. In the QMC model, quarks are treated as fundamental degrees of freedom and are confined inside the nucleon bag. These quarks confined inside the nucleons interact through the exchange of scalar fields $\sigma$ and $\delta$ and the vector fields $\omega$ and $\rho$. In the chiral SU(3) hadronic mean field model, hadrons are treated as point particle. As discussed before, the properties of nucleons in the chiral SU(3) hadronic mean field model are modified through the exchange of scalar mesons, $\sigma,\zeta$ and $\delta$ and vector mesons $\omega$ and $\rho$. In the QMC model, the effective masses of different mesons in the nuclear medium are calculated through the energy of static bag in which quark and antiquarks are confined. On the otherhand, as explained in detail in Sec. III, in chiral hadronic model, the interaction Lagrangian density, describing the interactions of mesons under study with the nucleons of the medium are written. The dispersion relations are obtained from the interaction Lagrangian to calculate the optical potentials of mesons in the nuclear medium. In Ref. [44], the possibilities of bound states for $K$, $D$ and $B$ mesons were explored whereas in our current work, bound states of $\eta,\eta^{\prime}$ and $\phi$ mesons with different nuclei were also considered.
We have not considered the effect of finite density and temperature of the nuclear medium on the binding energy of mesic nuclei. An estimate of this can be obtained considering following picture: consider the formation of mesic nuclei in a nuclear medium at finite temperature, the medium effects can be incorporated by replacing the vacuum mass of mesons with the in-medium value in the calculation of binding energy (appearing in definitions of reduced mass $\mu$ [in Eq. (22)] and the real and imaginary part of optical potentials [Eqs. (24) and 25)] at given density and temperature. As an example, consider the formation of $D^{0}$ mesic nuclei in the nuclear medium $\rho_{B}=\rho_{0}$ and $T=100$ MeV, at which the mass of $D^{0}$ meson is found to be $1799.57$ MeV. Using this value, the mass shift $\Delta m_{\psi}(\rho_{0})$ at center of ${}^{208}_{82}$b nuclei is found to be $-42.29$ MeV. This can be compared with the value $-64.98$ MeV which was calculated without considering the impact of nuclear medium on formation of bound states and mass of $D^{0}$ in free space was 1864 MeV (see Fig.4(a)). Using the modifications due to finite density and temperature, the values of binding energy for $1s$ and $1p$ states change to $-40.10$ and $-37.41$ MeV and can be compared values $-49.133$ and $-43.096$ MeV given in Table 2. Thus, we observe a reduction in the magnitude of binding energy at finite density and temperature of the medium. The decrease in the magnitude of binding energy is also expected for other cases where bound states with different mesons are formed due to attractive interactions in the medium. The many-body effects which may impact the density distributions of protons and neutrons inside the nuclei and may also cause additional attractive interaction and absorption [91] will be studied in future work. The present investigations on the formation of meson-nuclei bound states will be significant for the experiments of $\bar{\text{P}}$ at FAIR, WASA at COSY and J-PARC facilities.